OpenStudy (anonymous):

Choose one of the factors of x3 + 343. x - 7 x2 - 14x + 49 x2 + 7x + 49 None of the above

jimthompson5910 (jim_thompson5910):

hint: 343 = 7^3

jimthompson5910 (jim_thompson5910):

so we have a sum of cubes x^3 + 343 = x^3 + 7^3

OpenStudy (anonymous):

ur hot k bye

jimthompson5910 (jim_thompson5910):

in which you'll use the sum of cubes factoring formula a^3 + b^3 = (a+b)(a^2 - ab + b^2)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so its d

jimthompson5910 (jim_thompson5910):

what are 'a' and b in this case?

OpenStudy (anonymous):

@jim_thompson5910 a=x-7 b=x^2-14x+49...

OpenStudy (anonymous):

@jim_thompson5910